if a kid takes 4 tests and it is his average grade and he wants to score a 90 and his first 3 test grades were 84, 93, and 88 what would the last test score be to get an average of 90?: if a kid takes 4 tests and it is his average grade and he wants to score a 90 and his first 3 test grades were 84, 93, and 88 what would the last test score be to get an average of 90?
Answer by mangopeeler07(442)

Set up an equation for the average: (84+93+88+x)/4=90

Combine like terms:

(265+x)/4=90

Multiply both sides by 4.

265+x=360

Subtract 265 from both sides

x=95

So for the average grade to be 90, the person would have to get a 95.